Two particles of mass M and 2M are at a distance d apart, they are moving towards each other under mutual force of attraction. The acceleration of their centre of mass is_____________

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

F=G*2m*m/d2

acceleration of mass m= G*2m/d2

acceleration of mass 2m =-G*m/d2

so Acceleratin of center of mass ={m*G*2m/d2    -m*G*2m/d2 }/(2m+m)

                                               =0


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin